OpenStudy (anonymous):

5(c-8) -3(2c+12) =-84

OpenStudy (anonymous):

c=8

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[5(c-8) -3(2c+12) =-84\]\[5c-40 -6c+36 =-84\]

OpenStudy (anonymous):

because distributive property

OpenStudy (anonymous):

you distribute first.

OpenStudy (anonymous):

I somehow got 8/11

OpenStudy (anonymous):

Btw @joehutchinson6 answer is wrong.

OpenStudy (anonymous):

I know what I messed up

OpenStudy (anonymous):

\[5c−40−6c+36=−84\]combine like terms you get \[−4−c=−84\]

OpenStudy (anonymous):

does C = 80?

OpenStudy (anonymous):

you add 4 to both sides and you get -c =-80 multiply -1 to both sides and you are right c=80

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

dont forget to give me thumbs up!

OpenStudy (anonymous):

how sorry? just joined